# Asymptotically Unbiased Synthetic Control Methods by Distribution Matching

### v3

@@ 3 Recap of SCMs
This section reviews the original SCM proposed by Abadie and Gardeazabal (2003). To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ by using the weighted sum of $\{Y^{N}_{j,t}\}^{J}_{j=1}$ for $t\in\mathcal{T}_{1}$; that is, SCMs predict $Y^{N}_{0,t}$ as
@@
where $w_{0}\in\mathbb{R}$ and $w_{j}\in\mathbb{R}$ $(j\in\mathcal{J}^{U})$ are the weights. We refer to these weights as SC weights.
Several methods exist for estimating the SC weights. In the SCM proposed by Abadie and Gardeazabal (2003), SC weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1t},\dots,Y^{N}_{J,t}$. Let us define
$\Delta^{J}\coloneqq\{\bm{w}=(w_{1}\ w_{2}\ \dots\ w_{J})^{\top}\in[0,1]^{J}\mid\sum_{j\in\mathcal{J}^{U}}w_{j}=1\}.$
Then, Abadie and Gardeazabal (2003) estimates the SC weights as
$\bm{w}^{\mathrm{Abadie}}=\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t}\right)^{2},$
The counterfactual outcome is then predicted as
@@
Using the prediction $\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$, we estimate the treatment effect as $\widehat{\tau}^{\mathrm{Abadie}}_{t}=Y^{I}_{0,t}-\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$.
The literature on inference for SCMs has grown significantly, extending beyond the original proposals of Abadie and Gardeazabal (2003) and Abadie et al. (2010). Notable contributions include works by Li (2020), Shaikh and Toulis (2021), Cattaneo et al. (2021), and Chernozhukov et al. (2019). Doudchenko and Imbens (2016) proposes estimating the SC weight as
@@
where $\lambda_{1}>0$ and $\lambda_{2}>0$ are regularization parameters. The parameter space for $\widetilde{\bm{w}}$ can be restricted in various ways, such as setting $w_{0}=0$ and $\bm{w}=(w_{1}\ \cdots\ w_{J})^{\top}\in\Delta^{J}$.
4 Identification and Endogeneity
This section addresses the identification problem of treatment effects from the perspective of endogeneity.
@@ 4.1 Linear Models in Expected Outcomes
We reconsider the underlying assumptions in SCMs. Existing studies often posit some linear relationship between $Y^{N}_{0,t}$ and $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. One such relationship assumes the existence of $\bm{w}^{*}\in\Delta^{J}$ (Abadie et al., 2010):
@@
This assumption, known as the perfect pre-treatment fit (Ferman and Pinto, 2021), implies that the stochastic process of $Y^{N}_{0,1},\dots,Y^{N}_{0,T}$ can be completely replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$. In linear factor models, $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, and this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is considered unrealistic (note that this is a specific case and not a necessary condition for $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$).
A more natural assumption is that each term in the linear factor models of $Y^{N}_{0,t}$ can be replicated by the weighted sum of corresponding terms in the models of $Y^{N}_{0,1},\dots,Y^{N}_{0,T}$. A weaker assumption still is a linear relationship between $\mathbb{E}_{0,t}\left[Y^{N}_{0,t}\right]$ and $(\mathbb{E}_{j,t}\left[Y^{N}_{j,t}\right])_{j\in\mathcal{J}^{U}}$, which includes the replication of each term in linear factor models as a specific case.
@@
There exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$, it holds that
$\mathbb{E}_{0,t}\left[Y^{N}_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y^{N}_{j,t}\right],$ (2)
where $\Delta^{J}\coloneqq\{(w_{1},\dots,w_{J})\in\mathbb{R}^{J}\mid w_{j}\geq 0\ \forall j\ \mathrm{and}\ \sum_{j\in\mathcal{J}^{U}}w_{j}=1\}$.
Compared to Assumption 4.1, assuming $Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}Y_{j,t}$ is a stronger and less convincing assumption than (2) since $Y_{j,t}$ is a random variable. Therefore, in this study, we discuss SCMs based on (2). Ferman and Pinto (2021) explicitly considers Assumption 4.1.
@@ 4.2 Asymptotic Bias in Linear Factor Models
We consider the linear factor model as defined in Definition 2.1, along with Assumption 4.1. However, under the linear factor model, we adopt the following more specific assumption:
@@
Under the linear factor model from Definition 2.1, there exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$, the following conditions hold:
$c_{0}=\sum_{j\in[J]}w^{*}_{j}c_{j}\quad\text{and}\quad\mu_{0}=\sum_{j\in[J]}w^{*}_{j}\mu_{j}\ \ \forall t\in\mathcal{T},\ \forall j\in\mathcal{J}.$
Assumption 4.2 is a specific case of Assumption 4.1; that is, if Assumption 4.1 holds, then Assumption 4.2 also holds under the linear factor model defined in Definition 2.1.
Under the linear factor model with Assumption 4.2, Ferman and Pinto (2021) finds that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{t}$ is biased and not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero. We further make the following assumption for the linear factor models.
@@
For each $j\in\mathcal{J}$, the linear factor models 2.1 satisfy the following conditions as $T_{0}\to\infty$:
@@
Ferman and Pinto (2021) then demonstrates that the SCM proposed by Abadie and Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$ under these conditions.
@@
Under Assumptions 4.34.2, it holds that
@@
as $T_{0}\to\infty$, where $(c_{0},\mu_{0})\neq(\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j},\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j})$, unless $\sigma^{2}_{\epsilon}=0$ or $\widetilde{\Phi}\cap\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\{\bm{w}^{\top}\bm{w}\}\neq\emptyset$. Moreover, for $t\in\mathcal{T}_{1}$, it holds that
@@
To mitigate this issue, Ferman and Pinto (2021) proposes the demeaned SC estimator
@@
where $\overline{Y}^{N}_{j}=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}Y_{j,t}$. They estimate $\tau_{0,t}$ as $\widehat{\tau}^{\mathrm{FP}}_{0,t}=Y^{I}_{0,t}-\overline{Y}^{N}_{0}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{FP}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right)$. They show the following result.
@@ 4.3 Implicit Endogeneity
In this section, we interpret the asymptotic bias identified by Ferman and Pinto (2021) from the perspective of endogeneity.
To clarify this issue, we consider a simpler case using a least-squares estimator rather than $\bm{w}^{\mathrm{Abadie}}$:
@@
We demonstrate that $\widehat{\bm{w}}^{\mathrm{LS}}$ is asymptotically biased due to endogeneity. Define $\epsilon_{j,t}=Y_{j,t}-\mathbb{E}_{j,t}\left[Y_{j,t}\right]$. Assuming that (2) holds, the following linear regression model is applicable:
@@
Here, $Y_{0,t}=\mathbb{E}_{0,t}[Y_{0,t}]+\epsilon_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]+\epsilon_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}-\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\epsilon_{j,t}+\epsilon_{0,t}$. Note that $\mathbb{E}\left[\nu_{t}Y_{j,t}\right]\neq 0$, implying endogeneity due to a correlation between the regressor and the error term. Endogeneity is known to introduce bias in the least-squares estimator. Define $Q^{*}$ as a $(J\times J)$-matrix whose $(i,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\mathbb{E}_{i,t}[Y_{i,t}]\mathbb{E}_{j,t}[Y_{j,t}]$. Let $\Sigma$ be a $J\times J$ diagonal matrix whose $(j,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\sigma^{2}_{j,t}$.
@@
This theorem is derived from Section 5.6 of Greene (2003). The result implies that under the assumption of (2), there is an asymptotic bias in $\widehat{\bm{w}}^{\mathrm{LS}}$; that is, least-squares-type estimators do not converge to the true value $\bm{w}^{*}$.
This section highlights the identification of treatment effects and the problem of implicit endogeneity. Though the estimator $\bm{w}^{\mathrm{Abadie}}$ is widely used in existing studies, we find that it suffers from implicit endogeneity. This issue has been explicitly or implicitly discussed in previous studies (Chernozhukov et al., 2019; Ferman and Pinto, 2021; Ferman, 2021). Independently of our work, Fry (2024) explicitly addresses the bias as an endogeneity problem and applies the instrumental variable method to resolve the issue.
@@ 5 SCM by Density Matching
As discussed in Section 4, Assumption 4.1 is considered weak or incompatible for identification of $\bm{w}^{*}$ using least squares. Therefore, we consider adopting another assumption for identification. Specifically, we propose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$ and introduce a SCM based on the density-matching approach.
@@ 5.1 Identification and Estimation Strategy
Section 4 and the results in Ferman and Pinto (2021) imply that combining the assumption $\mathbb{E}_{0,t}\left[Y_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]$ with least squares estimators does not yield consistent estimators for $\bm{w}^{*}$. Hence, we consider another more realistic and theoretically favorable assumption.
We begin with the premise that the distribution of the target unit’s outcome, $Y^{N}_{0,t}$, is a linear combination of those of the control units’ outcomes, $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. For simplicity, in this subsection, we assume that $Y^{N}_{j,t}$ is a continuous random variable with a probability density function (PDF). We then consider a scenario where the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$:
@@
We discuss how to estimate SC weights that satisfy (3). One approach is to employ the optimal-transport argument, as used by Gunsilius (2023), which estimates SC weights by minimizing the Wasserstein distance. However, in this study, we provide a simpler algorithm for estimating SC weights since the method of Gunsilius (2023) requires a procedure that is more complex than the original SCMs.
Because the mixture model (3) holds, we establish the following equation:
$\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\quad\forall t\in\mathcal{T},\ \ \forall\xi\in\mathbb{R},$ (4)
where $\sqrt{-1}$ denotes the imaginary number. From this equation, we estimate $w^{*}$ using the following methodology. We consider the upper bound of the difference between the expected exponential transforms of $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{j,t}$:
$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\right|\mathrm{d}\xi$
$=\int_{\xi\in(-h,h)}\left|\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right)+\epsilon|\xi^{G}|\right|\mathrm{d}\xi$
$\leq\int_{\xi\in(-h,h)}\left(\sum_{\gamma\in\{1,2,\dots,G\}}\frac{|\sqrt{-1}^{\gamma}||\xi|^{\gamma}}{\gamma!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\epsilon|\xi^{G}|\right)\mathrm{d}\xi$
$=\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}.$
This formula implies that by minimizing $\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\Big{(}Y^{N}_{0,t}\Big{)}^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\Big{(}Y^{N}_{j,t}\Big{)}^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}$, we can accurately estimate $\bm{w}^{*}$ such that $\int_{\xi in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{0,t}\Big{)}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp\Big{(}\sqrt{-1}\xi Y^{N}_{j,t}\Big{)}\right]\right|\mathrm{d}\xi=0$. This estimation strategy can be interpreted as a variant of the method of moments, where we estimate SC weights to match all moments between the treated unit and the weighted sum of the untreated units.

### v4

@@ 3 Recap of SCMs
This section reviews the original SCM proposed by Abadie and Gardeazabal (2003). To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ using a weighted sum of $\{Y^{N}_{j,t}\}_{j=1}^{J}$ for $t\in\mathcal{T}_{1}$. Specifically, SCMs predict $Y^{N}_{0,t}$ as
@@
where $w_{0}\in\mathbb{R}$ and $w_{j}\in\mathbb{R}$ $(j\in\mathcal{J}^{U})$ are the SC weights.
Several methods exist for estimating the SC weights. In the SCM proposed by Abadie and Gardeazabal (2003), these weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. Define
$\Delta^{J}\coloneqq\left\{\bm{w}=(w_{1}\ w_{2}\ \dots\ w_{J})^{\top}\in[0,1]^{J}\mid\sum_{j\in\mathcal{J}^{U}}w_{j}=1\right\}.$
Then, Abadie and Gardeazabal (2003) estimate the SC weights as
$\bm{w}^{\mathrm{Abadie}}=\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\sum_{t\in\mathcal{T}_{0}}\left(Y_{0,t}-\sum_{j\in\mathcal{J}^{U}}w_{j}Y^{N}_{j,t}\right)^{2}.$
Recall that we referred to SCMs using this approach as the Least-Squares-SCMs (LS-SCMs). The counterfactual outcome is then predicted as
@@
Using $\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}$, we estimate the treatment effect as
$\widehat{\tau}^{\mathrm{Abadie}}_{t}=Y^{I}_{0,t}-\widehat{Y}^{N,\mathrm{Abadie}}_{0,t}.$
The literature on inference for SCMs has grown significantly, extending the original proposals by Abadie and Gardeazabal (2003) and Abadie et al. (2010). Notable contributions include works by Li (2020), Shaikh and Toulis (2021), Cattaneo et al. (2021), and Chernozhukov et al. (2019). Doudchenko and Imbens (2016) proposes estimating the SC weight as
@@
where $\lambda_{1}>0$ and $\lambda_{2}>0$ are regularization parameters.
4 Identification and Implicit Endogeneity
This section addresses the identification problem of treatment effects and points out the problem of implicit endogeneity in LS-SCMs.
@@ 4.1 Linear Models in Expected Outcomes
We revisit the assumptions underlying SCMs. In many existing studies, a linear relationship is assumed between $Y^{N}_{0,t}$ and $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. One possibility is positing the existence of $\bm{w}^{*}\in\Delta^{J}$ (Abadie et al., 2010) such that
@@
This assumption, often called the perfect pre-treatment fit (Ferman and Pinto, 2021), implies that the stochastic process $\left(Y^{N}_{0,t}\right)_{t\in\mathcal{T}_{0}}$ can be fully replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$ for $t\in\mathcal{T}_{0}$. In linear factor models where $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is often considered unrealistic. Note that this is a specific case and not necessary for $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$.
A more natural assumption is a linear model between $\mathbb{E}_{0,t}[Y^{N}_{0,t}]$ and $(\mathbb{E}_{j,t}[Y^{N}_{j,t}])_{j\in\mathcal{J}^{U}}$.
@@
There exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$,
$\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}].$ (2)
Compared to Assumption 4.1, the assumption $Y_{0,t}=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}Y_{j,t}$ is stronger and less plausible, given that $Y_{j,t}$ is random.
@@ 4.2 Asymptotic Bias in Linear Factor Models
We now consider the linear factor model introduced in Definition 2.1, along with Assumption 4.1. This is a specific case of (2). Under the linear factor model, we adopt the following more specialized assumption:
@@
Under the linear factor model from Definition 2.1, there exists a set $\widetilde{\Phi}\subset\Delta^{J}$ such that for all $\bm{w}^{*}=(w^{*}_{j})_{j\in\mathcal{J}^{U}}\in\widetilde{\Phi}$,
$c_{0}=\sum_{j\in[J]}w^{*}_{j}c_{j}\quad\text{and}\quad\mu_{0}=\sum_{j\in[J]}w^{*}_{j}\mu_{j}\quad\forall t\in\mathcal{T},\ \forall j\in\mathcal{J}.$
Assumption 4.2 is a specific case of Assumption 4.1; if Assumption 4.1 holds, then Assumption 4.2 also holds under the linear factor model in Definition 2.1.
Under these assumptions, Ferman and Pinto (2021) shows that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{0,t}$ suffers from bias and is not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero. We also adopt the following assumption regarding the linear factor model.
@@
For each $j\in\mathcal{J}$, the linear factor model in Definition 2.1 satisfies the following as $T_{0}\to\infty$:
@@
Under Assumptions 4.2 and 4.3, Ferman and Pinto (2021) concludes that the LS-SCM proposed by Abadie and Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$, as the following proposition.
@@
Under Assumptions 4.34.2, we have
@@
as $T_{0}\to\infty$, where
$(c_{0},\mu_{0})\neq\left(\sum_{j\in\mathcal{J}^{U}}c_{j}\overline{w}^{\mathrm{Abadie}}_{j},\sum_{j\in\mathcal{J}^{U}}\mu_{j}\overline{w}^{\mathrm{Abadie}}_{j}\right),$
unless $\sigma^{2}_{\epsilon}=0$ or $\widetilde{\Phi}\cap\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J}}\{\bm{w}^{\top}\bm{w}\}\neq\emptyset$. Furthermore, for $t\in\mathcal{T}_{1}$,
@@
To alleviate this bias problem, Ferman and Pinto (2021) proposes the demeaned SC estimator:
@@
where $\overline{Y}^{N}_{j}=\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}Y_{j,t}$. They then estimate $\tau_{0,t}$ as
$\widehat{\tau}^{\mathrm{FP}}_{0,t}=Y^{I}_{0,t}-\overline{Y}^{N}_{0}-\sum_{j\in\mathcal{J}^{U}}\widehat{w}^{\mathrm{FP}}_{j}\left(Y^{N}_{j,t}-\overline{Y}^{N}_{j}\right).$
They show that this approach can reduce the bias seen in the classical LS-SCM estimator. However, there still remains bias in the estimator, which does not vanish even if $T_{0}\to\infty$.
@@ 4.3 Implicit Endogeneity
We now interpret the asymptotic bias found by Ferman and Pinto (2021) as an endogeneity problem. We consider (2), that is, $\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}]$, which is more general than the linear factor model.
To illustrate, we consider a simpler case with the least-squares estimator rather than $\bm{w}^{\mathrm{Abadie}}$:
@@
We show that $\widehat{\bm{w}}^{\mathrm{LS}}$ is asymptotically biased due to endogeneity. Define $\epsilon_{j,t}=Y_{j,t}-\mathbb{E}_{j,t}[Y_{j,t}]$. Under (2), we obtain
@@
Noting that $\mathbb{E}[\nu_{t}Y_{j,t}]\neq 0$, we see that there is endogeneity, as the regressor and the error term are correlated. Endogeneity is known to induce bias in least-squares estimation. Let $Q^{*}$ be a $(J\times J)$ matrix whose $(i,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\mathbb{E}_{i,t}[Y_{i,t}]\mathbb{E}_{j,t}[Y_{j,t}]$, and let $\Sigma$ be a $J\times J$ diagonal matrix whose $(j,j)$-element is $\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\sigma^{2}_{j,t}$.
@@
This result follows from Section 5.6 of Greene (2003) and implies that, under (2), there is an asymptotic bias in $\widehat{\bm{w}}^{\mathrm{LS}}$; in other words, least-squares-type estimators do not converge to the true $\bm{w}^{*}$.
Thus, the identification of treatment effects faces a challenge due to implicit endogeneity. While the estimator $\bm{w}^{\mathrm{Abadie}}$ remains standard in many applications, it inherits this endogeneity issue. Such concerns have been noted in prior works (Chernozhukov et al., 2019; Ferman and Pinto, 2021; Ferman, 2021), and Fry (2024) independently frames the bias as an endogeneity problem, proposing an instrumental variable method to address it.
@@ 5 SCM by Density Matching
As discussed in Section 4, Assumption 4.1 can be weak or incompatible for identifying $\bm{w}^{*}$ when using least squares. We, therefore, adopt a different assumption for identification. Specifically, we propose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted combination of the densities of $Y^{N}_{j,t}$, and we introduce an SCM based on density matching.
@@ 5.1 Identification and Estimation Strategy
Section 4 and the results in Ferman and Pinto (2021) indicate that combining
$\mathbb{E}_{0,t}\left[Y_{0,t}\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[Y_{j,t}\right]$
with least squares estimators do not yield consistent estimators of $\bm{w}^{*}$. We thus adopt mixture models as a more theoretically favorable but still realistic assumption.
We begin by positing that the distribution of the target unit’s outcome $Y^{N}_{0,t}$ can be expressed as a linear combination of the distributions of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. For simplicity, in this subsection, we assume that $Y^{N}_{j,t}$ is a continuous random variable with a probability density function (PDF). Under this scenario, we suppose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted sum of the densities of $Y^{N}_{j,t}$:
@@
We next discuss estimating SC weights that satisfy (3). One approach employs optimal transport to minimize the Wasserstein distance (Gunsilius, 2023). In this study, we propose a simpler estimation procedure, as the approach of Gunsilius (2023) involves more complexity than the original SCMs.
Because the mixture model (3) holds, we obtain:
$\mathbb{E}_{0,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{0,t}\right)\right]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}\left[\exp\left(\sqrt{-1}\xi Y^{N}_{j,t}\right)\right]\quad\forall t\in\mathcal{T},\ \forall\xi\in\mathbb{R},$
where $\sqrt{-1}$ denotes the imaginary unit. From this, we can estimate $\bm{w}^{*}$ by considering the upper bound of the difference between the expected exponential transforms of $Y^{N}_{0,t}$ and the corresponding weighted sum for $Y^{N}_{j,t}$. Specifically, we have
$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{0,t})\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{j,t})\right]\right|\mathrm{d}\xi$
$=\int_{\xi\in(-h,h)}\Biggl{|}\sum_{\gamma\in\{0,1,2,\dots,G\}}\frac{\sqrt{-1}^{\gamma}\xi^{\gamma}}{\gamma!}\left(\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right)+\epsilon\left|\xi^{G}\right|\Biggr{|}\mathrm{d}\xi$
$\leq\int_{\xi\in(-h,h)}\Biggl{(}\sum_{\gamma\in\{1,2,\dots,G\}}\frac{|\sqrt{-1}^{\gamma}||\xi|^{\gamma}}{\gamma!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|+\epsilon|\xi^{G}|\Biggr{)}\mathrm{d}\xi$
$=\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\mathbb{E}_{0,t}\left[\left(Y^{N}_{0,t}\right)^{\gamma}\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\left(Y^{N}_{j,t}\right)^{\gamma}\right]\right|+\frac{2\epsilon}{G+1}h^{G+1}.$
Minimizing
$\sum_{\gamma\in\{1,2,\dots,G\}}\frac{2h^{\gamma+1}}{(\gamma+1)!}\left|\frac{1}{T_{0}}\sum_{t\in\mathcal{T}_{0}}\left(\left(Y^{N}_{0,t}\right)^{\gamma}-\sum_{j\in\mathcal{J}^{U}}w_{j}\left(Y^{N}_{j,t}\right)^{\gamma}\right)\right|+\frac{2\epsilon}{G+1}h^{G+1}$
thus yields an accurate estimate of $\bm{w}^{*}$ that enforces
$\int_{\xi\in(-h,h)}\left|\mathbb{E}_{0,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{0,t})\right]-\sum_{j\in\mathcal{J}^{U}}w_{j}\mathbb{E}_{j,t}\left[\exp(\sqrt{-1}\xi Y^{N}_{j,t})\right]\right|\mathrm{d}\xi=0.$
This approach can be seen as a variant of the method of moments, where we match all moments of $Y^{N}_{0,t}$ with those of the weighted sum of the untreated units.